OpenStudy (anonymous):

In the right triangles below, DR = 16 feet, DC = 12 feet, and BF = 18 feet. What is QF?

OpenStudy (anonymous):

i need the picture lol

OpenStudy (anonymous):

set up a ratio: BF/QF = DC/DR 18/x = 12/16 18*16 = 12x 18*16/12 = x 288/12 = x 24 = x